What is the magnitude of the acceleration of a proton in a uniform electric field of magnitude 7197 N/C? Give your answer in Standard SI units. Remember that when entering a number into the computer using scientific notation you must enter it using the "e" notation (for example 6.02 x 1023 must be entered as 6.02e23.)
Answer #1

Dear student. feel Free to ask doubts in comment. I'll surely reply within few seconds.

E = 7197 N/C Force = F = qe - acceleration = F =106 x 10 x 7197 1-67 X10-27 = 16-8953x10) m/s2 = 16.8953 e 117 m/s2
Hope you give thumbs up.

Thanks!